OpenStudy (anonymous):

Choose the ratio that is equivalent to 1/7 opts:3/21 14/2 4/11 3/14

OpenStudy (saifoo.khan):

A,

OpenStudy (saifoo.khan):

multiply 1/3 by 3

OpenStudy (callisto):

3/21 3/21 = (1x3) / (7x3) = (1/7)(3/3) = (1/7)x1 = 1/7